For high-pass active filter with corner

frequency 8 kHz, then R and C are:


Select one:

a. 2.5 k Ω , 7.9557 nF

b. 2 k Ω , 7.9557 nF

c. 5 k Ω , 7:9557 nF nF

d. 10 k Ω , 7.9557 nF

e. 1 k Ω ,7.9557 nF


1
Expert's answer
2021-06-07T05:46:03-0400

Corner frequency ,


"f_c= \frac{1}{2\pi RC}"


From option C= 7.9557 nF is selected

Corner frequency = 8kHz


Hence Resistance,

"R=\frac{1}{2\pi f_c C}= \frac{1}{2\pi\times 8\times10^3 \times7.9557\times10^{-9}}"


"R=2500 \varOmega =2.5k\varOmega"


Hence option A is the correct answer